OpenStudy (liv1234):

Find the least common multiple of x^2+x-12 and x^2+2x-15. A)(x + 4)(x - 3)(x + 5) B)(x - 3)(x + 5)(x - 4) C)(x - 4)(x - 3)(x - 5) D)(x + 4)(x - 5)(x - 3) @skullpatrol @pooja195

OpenStudy (liv1234):

Oh wait, also I got the answer of (x-3) (x+4) (x+5), but I don't know which option to choose :/

OpenStudy (abdullahm):

Factor x^2 + 2x -15 and x^2 + x - 12. What do you get when you factor it?

OpenStudy (liv1234):

I feel like I should choose option A, but it doesn't have the same order as the answer that I got.

OpenStudy (abdullahm):

If that's what you got, then your answer is A! Since (a)(b)(c) = (b)(a)(c) =(b)(c)(a) =(c)(b)(a) = (c)(a)(b) It doesn't mater which order the terms are written in when you are multiplying them!

OpenStudy (liv1234):

It doesn't? Ohh okay, because I know that a lot of times you have to make sure terms are in a certain order, or the answer isn't correct.

OpenStudy (liv1234):

It was correct! Thank you for letting me know that, now I know that when I come across a question like this, (if it is multiplication), it doesn't matter what order they are in as long as they have the right terms.

OpenStudy (abdullahm):

When you're multiplying the terms, order doesn't matter since you will get the same value at the end. The same with if you're just adding the terms.

OpenStudy (abdullahm):

This is related to the associative property and commutative property. Be careful when you're subtracting and dividing, however!

OpenStudy (abdullahm):

It was my pleasure! :)